f(x)=4x-7, find f(x+h)-f(x)/h
I believe I am supposed to pluf the number given into the equation.

Answer by Alan3354(69443)

f(x)=4x-7, find f(x+h)-f(x)/h
--------
f(x+h) = 4(x+h)-7
f(x+h)-f(x) = 4(x+h)-7 -(4x-7) = 4x+4h - 4x = 4h
4h/h = 4